Reports 1-1 of 1 Clear search Modify search
Injection system (Laser power stabilization)
derossi, chiummo, gosselin, spinicelli - 17:44 Wednesday 22 September 2021 (53228) Print this report
recovering some power from PSTAB offset

In order to reach 40 W injected into the ITF, as agreed in the PSL/INJ meeting, we changed the correction offset voltage of the AOM of the power stabilization loop.

We proceeded as follows:

  • we opened PSTAB loop
  • we unplugged the corrections (named out corr) at 14:13:50 UTC --> the power at the output of the PMC increased by 5%
  • the potentiometer for the corrections was set to 8.5. We set it 0 and plugged again the corrections at 14:15:28 UTC
  • we increased the corrections voltage step by step from 14:17:00 UTC to 14:25:00 UTC (1st plot)
  • we verified from a trend that the minimum correction range needed is 0.85 Vpp (2nd plot)
  • we set the offset on the potentiometer to 2.5 (14:52:22 UTC), to be in the range where the voltage effect is known. The 3rd plot shows the corrections and the error signal before and after this adjustmnet.
Images attached to this report
Comments to this report:
chiummo - 9:52 Thursday 23 September 2021 (53239) Print this report

The power injected to the ITF is around 40W after the overall settling overnight. It also seems that the PMC throughput has slightly improved, but it is difficult to assess because noise on the recorded reflected power PMC_REFL is such to hide the expected increase.

Images attached to this comment
Search Help
×

Warning

×